OpenStudy (j2lie):

How do I turn X^2 - 7 = 4x + 5 into completing the square problem?

OpenStudy (anonymous):

\[(x-2)^{2}=8\]

OpenStudy (j2lie):

How did you get the eight?

Nnesha (nnesha):

first you have to move 4x on left side

OpenStudy (dumbcow):

\[x^2 -4x = 12\] \[x^2 -4x + (\frac{4}{2})^2 = 12 + (\frac{4}{2})^2\]

Nnesha (nnesha):



Nnesha (nnesha):

then move 7 on right side

Nnesha (nnesha):